# Law of Large Numbers for continuous $N$-particle ensembles at fixed temperature

### v1

@@ 1.4 Methodology
Theorem 3.4 has two directions, the “if” and “only if” parts; we take quite different approaches for their proofs. In the “if” direction, the key is that we are able to extract the moments from the Bessel generating functions by means of the Dunkl operators. While Dunkl operators were introduced in a purely algebraic background and historically play a role in the study of Calogero–Moser–Sutherland quantum many-body system, they have recently found surprising applications in random matrix theory and statistical mechanics, see e.g. [BCG22, Xu23, KX24, GXZ24, Yao25].
One new feature in our argument, compared to the previous literature, is that we consider the expansions of logarithms of Bessel generating functions in the basis of power sum symmetric polynomials instead of monomial symmetric polynomials. While the coefficients for the latter basis can be more easily extracted by taking partial derivatives, hence seeming like a more natural choice, their asymptotics do not give the right characterization of the LLN.
In the “only if” direction, while it seems doable to reverse the arguments developed in the Dunkl actions, we took a different tool from algebraic combinatorics, which greatly simplifies the proof. This is a new topological expansion, discovered by Chapuy–Dolega [CD22], of the $\theta$-generalized tau-function in terms of a class of combinatorial objects called constellations, which are certain maps on closed orientable and non-orientable surfaces. We take a specialization of this formula and identify the tau-function with the multivariate Bessel function; then the topological expansion highlights clearly what the asymptotically relevant terms are. It is worth noting that there have been several connections between topological expansions and random matrices; see, e.g. [BG13, CE06, BGF20]. Although topological expansions appeared in very different ways compared to our approach, we believe that there is a deeper structural explanation lying behind its connection with random matrices. We hope to investigate such connections in future work.
@@ 2.1 Conventions
In this paper, $N$ denotes a positive integer that will tend to infinity. For a sequence of infinite sums $f_{N}=\sum_{j=-\infty}^{n}a_{j}(N)N^{j}$, $a_{j}(N)\rightarrow a_{j}$ as $N\to\infty$ for all $j$’s, we denote the $N$-degree of $\{f_{N}\}$ by the minimal integer $n$, such that $a_{j}=0$ for all $j\geq n+1$.
We will work with $N$ variables $x_{1},\dots,x_{N}$ and let $\vec{x}=(x_{1},\dots,x_{N})$. The partial derivatives will be denoted
@@ 2.2 Dunkl operators
They pairwise commute, $\mathcal{D}_{i}^{N,\theta}\mathcal{D}_{j}^{N,\theta}=\mathcal{D}_{j}^{N,\theta}\mathcal{D}_{i}^{N,\theta}$, for all $i,j$, and their symmetric joint eigenfunctions, discussed in the next subsection, will play an important role in this paper. We can view $\mathcal{D}_{i}^{N,\theta}$ as operators acting on holomorphic functions on domains of $\mathbb{C}^{N}$ that are symmetric with respect to permutations of the coordinates.
Further, we will use the symmetric operators
$\mathcal{P}_{k}^{N,\theta}:=\sum_{i=1}^{N}{\Big(\mathcal{D}_{i}^{N,\theta}\Big)^{k}},\qquad k=1,2,\dots.$
For simplicity, we drop the superscripts $N,\theta$, meaning that $\mathcal{D}_{i}^{N,\theta}$ and $\mathcal{P}_{k}^{N,\theta}$ will be simply denoted by $\mathcal{D}_{i}$ and $\mathcal{P}_{k}$.
@@ 2.3 Multivariate Bessel functions
Let $a_{1},\dots,a_{N}\in\mathbb{C}$ be arbitrary. There exists a unique holomorphic function $F(x_{1},\dots,x_{N})$ that is symmetric with respect to its variables, normalized by $F(0^{N})=1$, and solves the differential equations
@@
If we denote this solution by $B_{(a_{1},\dots,a_{N})}(x_{1},\dots,x_{N};\theta)$, then the correspondence $(a,x)\mapsto B_{a}(x;\theta)$ admits an extension to an entire function on $2N$ variables.
Due to the uniqueness of Theorem 2.1, it follows that $B_{(a_{1},\dots,a_{N})}(x_{1},\dots,x_{N};\theta)$ is also symmetric with respect to the variables $a_{i}$. When $a_{1}<\dots<a_{N}$, the multivariate Bessel function admits the following integral representation, see [Cue21, Proposition 3.3]:
@@
where $\Lambda_{\theta}(\vec{y},a)$ is the $\theta$-corners process defined in Eqn. (5), the $\frac{N(N-1)}{2}$–dimensional integral is over $k$-tuples $y^{k}=(y^{k}_{1},\dots,y^{k}_{k})$, $k=1,2,\dots,N-1$, and the conditions $y^{i}\prec y^{i+1}$ mean that $y^{i+1}_{1}\leq y^{i}_{1}\leq y^{i+1}_{2}\leq\dots y^{i+1}_{i}\leq y^{i}_{i}\leq y^{i+1}_{i+1}$.
@@ 2.4 Bessel generating functions
Following [BCG22], let $\mathcal{M}_{N}$ be the set of Borel probability measures on ordered real $N$-tuples $a_{1}\leq a_{2}\leq\dots\leq a_{N}$ and, for any $R>0$, let $\mathcal{M}_{N}^{R}\subset\mathcal{M}_{N}$ be the subset of measures $\mu$ such that
$\int_{a_{1}\leq a_{2}\leq\dots\leq a_{N}}e^{NR\max_{i}|a_{i}|}\,\mu(\textrm{d}a_{1},\dots,\textrm{d}a_{N})<\infty.$ (9)
Proposition-Definition 2.2.
If $\mu\in\mathcal{M}_{N}^{R}$, for some $R>0$, then the integral
$G_{\theta}(x_{1},\dots,x_{N};\mu):=\int_{a_{1}\leq a_{2}\leq\dots\leq a_{N}}B_{(a_{1},\dots,a_{N})}(x_{1},\dots,x_{N};\theta)\mu(\textrm{d}a_{1},\dots,\textrm{d}a_{N})$
@@
that will be called the Bessel generating function of $\mu$.
The measures in $\mathcal{M}_{N}^{R}$ will be called $R$-exponentially decaying. We also let $\mathcal{M}_{N}^{>0}:=\bigcup_{R>0}{\mathcal{M}_{N}^{R}}$, and all measures in $\mathcal{M}_{N}^{>0}$ will be called exponentially decaying. Any exponentially decaying measure $\mu$ admits a Bessel generating function $G_{\theta}(x_{1},\dots,x_{N};\mu)$ that is a holomorphic function in a neighborhood of the origin. By the normalization and symmetry of $B_{(a_{1},\dots,a_{N})}(x_{1},\dots,x_{N};\theta)$, it follows that
@@
where $a_{1}\leq a_{2}\leq\dots\leq a_{N}$ is $\mu$-distributed in the right hand side of the equality.
@@ 3 Statement of the main theorem
For the statement of the main theorem, let $\{\mu_{N}\in\mathcal{M}_{N}\}_{N\geq 1}$ be a sequence of exponentially decaying probability measures on real $N$-tuples $a=(a_{1}<\cdots<a_{N})$. The random variables of interest are
$p_{k}^{N}:=\sum_{i=1}^{N}{\left(\frac{a_{i}}{N}\right)^{k}},\qquad k\geq 1.$
@@
We say that $\{\mu_{N}\in\mathcal{M}_{N}\}_{N\geq 1}$ satisfies the fixed temperature LLN if there exists a sequence of real numbers $\{m_{k}\}_{k\geq 1}$, such that
@@
for all integers $s\geq 1$ and $k_{1},\dots,k_{s}\geq 1$.
Let
@@
be the Bessel generating function of $\mu_{N}$, for all $N\geq 1$. Further, assume that $\theta>0$ is a fixed constant, independent of $N$. Since $G_{N,\theta}(0^{N})=1$, then $\ln{G_{N,\theta}(\vec{x})}$ defines a holomorphic function in a neighborhood of $(0^{N})\in\mathbb{C}^{N}$. In particular, this means that each $\ln{G_{N,\theta}(\vec{x})}$ has a Taylor expansion around $(0^{N})$ of the form
@@
where the sum is over partitions $\nu$ with length $\ell(\nu)\leq N$, and $M_{\nu}(\vec{x})$ are the monomial symmetric polynomials. The Taylor coefficients $b_{\nu}^{N}$ can be calculated as
@@
To state the main theorem, we actually need to expand $\ln{G_{N,\theta}(\vec{x})}$ in the basis of power sums $\{p_{\lambda}(\vec{x})\}_{\lambda\colon\ell(\lambda)\leq N}$, where
@@
Since one can write $M_{\nu}(\vec{x})$ as a unique linear combination of power sums if and only if $|\nu|\leq N$, such expansion will be well-defined only up to degree $N$.
@@
We say that the sequence $\{\mu_{N}\in\mathcal{M}_{N}\}_{N\geq 1}$ is $\theta$-LLN-appropriate if the corresponding Bessel generating functions $G_{N,\theta}$ are such that if we expand, up to degree $N$,
@@
then there exist a sequence of real numbers $\{\kappa_{d}\}_{d\geq 1}$ such that
@@
    $\lim_{N\to\infty}{\frac{a_{(d)}^{N}}{N}}=\frac{\theta^{1-d}\kappa_{d}}{d}$, for all integers $d\geq 1$,
@@
    $\lim_{N\to\infty}{\frac{a_{\lambda}^{N}}{N^{\ell(\lambda)}}}=0$, for all partitions $\lambda$ with $\ell(\lambda)\geq 2$.
@@
Let $k\geq 1$ be an arbitrary integer. A Lukasiewicz path of length $k$ is a lattice path $\Gamma$ that starts at $(0,0)$, ends at $(k,0)$, has exclusively steps of the form $(1,d)$, for some d=1,0,1,2,d=-1,0,1,2,..., and stays in the first quadrant. We denote the set of all Lukasiewicz paths of length $k$ by $\mathbf{L}(k)$.
@@
Let $\theta>0$ be fixed. The sequence $\{\mu_{N}\}_{N\geq 1}$ satisfies the fixed temperature LLN (according to Definition 3.1) if and only if it is $\theta$-LLN-appropriate (according to Definition 3.2). In that case, the sequences $\{m_{k}\}_{k\geq 1}$ and $\{\kappa_{d}\}_{d\geq 1}$ from Definitions 3.1 and 3.2 uniquely determined each other, according to the equalities:
$m_{k}=\sum_{\Gamma\in\mathbf{L}(k)}{\prod_{d\geq 1}{(\kappa_{d})^{\#\text{steps $(1,d-1)$ of }\Gamma}}},\quad k\geq 1.$ (13)
We note that the coefficients $a_{\lambda}^{N}$ in the expansion (12) can be expressed in terms of $\ln{G_{N,\theta}(\vec{x})}$. This follows from the explicit expression for the monomial symmetric polynomials $M_{\mu}(\vec{x})$ in terms of the power-sum basis $p_{\lambda}(\vec{x})$ (see, for example, [Zab, Proposition 2.13]), by comparing the expansions (10) and (12), and using (11). Explicitly,
@@
In this way, conditions (a)-(b) in Definition 3.2 can be rewritten in terms of the Taylor coefficients of $\ln{G_{N,\theta}(\vec{x})}$. As a result, Theorem 3.4 answers the question posed in [BCG22], of finding necessary and sufficient conditions for a sequence $\{\mu_{N}\}_{N\geq 1}$ to satisfy the fixed temperature LLN in terms of the corresponding Bessel generating functions. Moreover, equation (14) coincides with the preliminary calculations in that appendix. For example, for the two partitions of size $2$, we have
@@
Let a(N)=(a,,a)a(N)=(a,...,a) where $a$ is a Gaussian random variable with mean 0 and variance $N$. Since $B_{(a,...,a)}(\vec{x};\theta)=\exp\left(a\sum_{i=1}^{N}x_{i}\right)$ (by taking a limit of Eqn. (8)), we have
$G_{N,\theta}(\vec{x})=\int_{-\infty}^{\infty}{\frac{1}{\sqrt{2\pi N}}\exp^{-\frac{a^{2}}{2N}}\exp\left(a\sum_{i=1}^{N}x_{i}\right)da}\\ =\exp\left(\frac{N}{2}p_{1}(\vec{x})^{2}\right)=\exp\left(\frac{N}{2}M_{(2)}(\vec{x})+NM_{(1,1)}(\vec{x})\right).$
One can check that $a(N)$ satisfies the fixed temperature LLN with $m_{k}=0$ for all $k$’s. This is revealed by the fact that $\lim_{N\rightarrow\infty}\frac{a^{N}_{(1,1)}}{N^{2}}=0$ and $a^{N}_{\lambda}=0$ for all $\lambda\neq(1,1)$. On the other hand, $G_{N,\theta}(\vec{x})$ does not satisfy the conditions proposed by [BCG22, Claim 9.1], which are given by monomials of partial derivatives.

### v2

@@ 1.4 Methodology
Theorem 3.4 has two directions, the “if” and “only if” parts; we take quite different approaches for their proofs. In the “if” direction, the key is that we are able to extract the moments from the Bessel generating functions by means of the Dunkl operators. While Dunkl operators were introduced in a purely algebraic background and historically play a role in the study of Calogero–Moser–Sutherland quantum many-body system, they have recently found many applications in random matrix theory and statistical mechanics; see [BCG22, Xu23, KX24, GXZ24, Yao25].
One new feature in our argument, compared to the previous literature, is that we consider the expansions of logarithms of Bessel generating functions in the basis of power sum symmetric polynomials instead of monomial symmetric polynomials. While the coefficients for the latter basis can be more easily extracted by taking partial derivatives, hence appearing like a more natural choice, their asymptotics do not give the right characterization of the LLN.
In the “only if” direction, while it seems doable to reverse the arguments developed in the forward direction, we employed a different tool that greatly simplified the proof. This is the new topological expansion, due to Chapuy–Dolega [CD22], of the $\theta$-generalized tau-function in terms of the combinatorial objects called constellations, which are certain maps on closed orientable and non-orientable surfaces. We take a specialization of this formula and identify the tau-function with the multivariate Bessel function; then the topological expansion highlights what the asymptotically relevant terms are. It is worth noting that there have been other connections between topological expansions and random matrices; see, e.g. [BG13, CE06, BGF20]. Although topological expansions arise in very different ways compared to our approach, there might be a deeper structural explanation lying behind their connection to random matrices; we hope to investigate this in future works.
@@ 2.1 Conventions
In this paper, $N$ denotes a positive integer that will tend to infinity. We will work with $N$ variables $x_{1},\dots,x_{N}$ and let $\vec{x}=(x_{1},\dots,x_{N})$. The partial derivatives will be denoted
@@ 2.2 Dunkl operators
They pairwise commute, $\mathcal{D}_{i}^{N,\theta}\mathcal{D}_{j}^{N,\theta}=\mathcal{D}_{j}^{N,\theta}\mathcal{D}_{i}^{N,\theta}$, for all $i,j$, and their symmetric joint eigenfunctions, discussed in the next subsection, will play an important role in this paper. We can view $\mathcal{D}_{i}^{N,\theta}$ as operators acting on holomorphic functions on domains of $\mathbb{C}^{N}$ that are symmetric with respect to permutations of the coordinates. We will also use the symmetric operators
$\mathcal{P}_{k}^{N,\theta}:=\sum_{i=1}^{N}{\Big(\mathcal{D}_{i}^{N,\theta}\Big)^{k}},\qquad k=1,2,\cdots.$
For simplicity, we drop the superscripts $N,\theta$, so that $\mathcal{D}_{i}^{N,\theta},\mathcal{P}_{k}^{N,\theta}$ will simply be denoted $\mathcal{D}_{i},\mathcal{P}_{k}$.
@@ 2.3 Multivariate Bessel functions
Let $a_{1},\dots,a_{N}\in\mathbb{C}$ be arbitrary. There exists a unique holomorphic function $F(x_{1},\dots,x_{N})$, normalized by $F(0^{N})=1$, that is symmetric with respect to its variables and solves the differential equations
@@
If we denote this solution by $B_{(a_{1},\dots,a_{N})}(x_{1},\dots,x_{N};\theta)$, then the map $(a,x)\mapsto B_{a}(x;\theta)$ can be extended to an entire function on $2N$ variables.
Due to the uniqueness of Theorem 2.1, it follows that $B_{(a_{1},\dots,a_{N})}(x_{1},\dots,x_{N};\theta)$ is also symmetric with respect to the variables $a_{i}$. When $a_{1}<\dots<a_{N}$, the multivariate Bessel function admits the following integral representation (see [Cue21, Prop. 3.3]):
@@
where $\Lambda_{\theta}(\vec{y},a)$ is the $\theta$-corners process defined in Eqn. (5), the $\frac{N(N-1)}{2}$–dimensional integral is over $k$-tuples $y^{k}=(y^{k}_{1},\dots,y^{k}_{k})$, for $k=1,2,\dots,N-1$, and the conditions $y^{i}\prec y^{i+1}$ mean that $y^{i+1}_{1}\leq y^{i}_{1}\leq y^{i+1}_{2}\leq\dots y^{i+1}_{i}\leq y^{i}_{i}\leq y^{i+1}_{i+1}$.
@@ 2.4 Bessel generating functions
Following [BCG22], let $\mathcal{M}_{N}$ be the set of Borel probability measures on ordered real $N$-tuples $a_{1}\leq\dots\leq a_{N}$ and, for any $R>0$, let $\mathcal{M}_{N}^{R}\subset\mathcal{M}_{N}$ be the subset of measures $\mu$ such that
$\int_{a_{1}\leq\dots\leq a_{N}}e^{NR\max_{i}|a_{i}|}\,\mu(\textrm{d}a_{1},\dots,\textrm{d}a_{N})<\infty.$ (9)
Proposition-Definition 2.2 (Lemma 2.9 of [BCG22]).
If $\mu\in\mathcal{M}_{N}^{R}$ for some $R>0$, then the integral
$G_{\theta}(x_{1},\dots,x_{N};\mu):=\int_{a_{1}\leq\dots\leq a_{N}}B_{(a_{1},\dots,a_{N})}(x_{1},\dots,x_{N};\theta)\mu(\textrm{d}a_{1},\dots,\textrm{d}a_{N})$
@@
that is called the Bessel generating function of $\mu$.
The measures in $\mathcal{M}_{N}^{R}$ will be called $R$-exponentially decaying. We also let $\mathcal{M}_{N}^{>0}:=\bigcup_{R>0}{\mathcal{M}_{N}^{R}}$, and call all measures in $\mathcal{M}_{N}^{>0}$ exponentially decaying. Any $\mu\in\mathcal{M}_{N}^{>0}$ admits a Bessel generating function $G_{\theta}(x_{1},\dots,x_{N};\mu)$ that is a holomorphic function in a neighborhood of the origin. By the normalization and symmetry of $B_{(a_{1},\dots,a_{N})}(x_{1},\dots,x_{N};\theta)$, it follows that
@@
where, in the right hand side, $a_{1}\leq\dots\leq a_{N}$ is $\mu$-distributed.
@@ 3 Statement of the main theorem
Let $\{\mu_{N}\in\mathcal{M}_{N}\}_{N\geq 1}$ be a sequence of measures on real $N$-tuples $a=(a_{1}<\cdots<a_{N})$. The random variables of interest are
$p_{k}^{N}:=\sum_{i=1}^{N}{\left(\frac{a_{i}}{N}\right)^{k}},\qquad k=1,2,\cdots.$
@@
We say that $\{\mu_{N}\in\mathcal{M}_{N}\}_{N\geq 1}$ satisfies the fixed temperature LLN if there exists a sequence of real numbers $\{m_{k}\}_{k\geq 1}$ such that
@@
for all integers $s\geq 1$, and $k_{1},\dots,k_{s}\geq 1$.
Assume that $\mu_{N}\in\mathcal{M}_{N}^{>0}$ and let
@@
be the Bessel generating function of $\mu_{N}$, for all $N\geq 1$. Since $G_{N,\theta}(0^{N})=1$, then $\ln{G_{N,\theta}(\vec{x})}$ defines a holomorphic function in a neighborhood of $(0^{N})\in\mathbb{C}^{N}$. In particular, this means that each $\ln{G_{N,\theta}(\vec{x})}$ has a Taylor expansion around $(0^{N})$ of the form
@@
where the sum is over partitions $\nu$ of length $\ell(\nu)\leq N$, and $M_{\nu}(\vec{x})$ are the monomial symmetric polynomials. The Taylor coefficients $b_{\nu}^{N}$ can be calculated as
@@
To state the main theorem, we need to expand $\ln{G_{N,\theta}(\vec{x})}$ in the power sum basis $\{p_{\lambda}(\vec{x})\}_{\lambda\colon\ell(\lambda)\leq N}$, where
@@
Since one can write $M_{\nu}(\vec{x})$ as a unique linear combination of power sums if and only if $|\nu|\leq N$, such expansion is well-defined only up to degree $N$.
@@
The sequence $\{\mu_{N}\in\mathcal{M}_{N}^{>0}\}_{N\geq 1}$ is $\theta$-LLN-appropriate if the corresponding Bessel generating functions $G_{N,\theta}$ are such that, if we expand $\ln{G_{N,\theta}(\vec{x})}$ up to degree $N$,
@@
then there exists a sequence of real numbers $\{\kappa_{d}\}_{d\geq 1}$ such that
@@
    $\lim_{N\to\infty}{\frac{a_{(d)}^{N}}{N}}=\frac{\theta^{1-d}}{d}\cdot\kappa_{d}$, for all $d\in\mathbb{Z}_{\geq 1}$,
@@
    $\lim_{N\to\infty}{\frac{a_{\lambda}^{N}}{N^{\ell(\lambda)}}}=0$, whenever $\ell(\lambda)\geq 2$.
@@
Let $k\geq 1$ be an arbitrary integer. A Łukasiewicz path of length $k$ is a lattice path $\Gamma$ that starts at $(0,0)$, ends at $(k,0)$, has steps of the form $(1,d)$, for some $d\in\{-1,0,1,2,\dots\}$, and stays in the first quadrant, i.e. does not contain vertices with negative $y$-coordinates. We denote the set of all Łukasiewicz paths of length $k$ by $\mathbf{L}(k)$.
@@
Let $\theta>0$ be fixed. The sequence $\{\mu_{N}\}_{N\geq 1}$ satisfies the fixed temperature LLN (according to Definition 3.1) if and only if it is $\theta$-LLN-appropriate (according to Definition 3.2). In that case, the sequences $\{m_{k}\}_{k\geq 1}$ and $\{\kappa_{d}\}_{d\geq 1}$ from Definitions 3.1 and 3.2 are uniquely determined by each other, according to the equalities:
$m_{k}=\sum_{\Gamma\in\mathbf{L}(k)}{\prod_{d\geq 1}{(\kappa_{d})^{\#\text{steps $(1,d-1)$ of }\Gamma}}},\text{ for all }k\geq 1.$ (13)
The coefficients $a_{\lambda}^{N}$ in the expansion (12) can be expressed in terms of $\ln{G_{N,\theta}(\vec{x})}$. This follows from the explicit expression for the monomial symmetric polynomials $M_{\mu}(\vec{x})$ in terms of the power-sum basis $p_{\lambda}(\vec{x})$ (see, e.g., [Zab, Prop. 2.13]), as well as equations (10), (11) and (12). Explicitly,
@@
In this way, conditions (a)-(b) in Definition 3.2 can be rewritten in terms of the Taylor coefficients of $\ln{G_{N,\theta}(\vec{x})}$. As a result, Theorem 3.4 answers the question posed in [BCG22] of finding necessary and sufficient conditions for a sequence $\{\mu_{N}\}_{N\geq 1}$ to satisfy the fixed temperature LLN in terms of the corresponding Bessel generating functions. Moreover, equation (14) coincides with the preliminary calculations in that appendix. For example, for the two partitions of size $2$, we have
@@
Let $a(N)=(a,\dots,a)$, for all $N\geq 1$, where $a$ is a Gaussian random variable of mean 0 and variance $N$. Since $B_{(a,\dots,a)}(\vec{x};\theta)=\exp\left(a\sum_{i=1}^{N}{x_{i}}\right)$ (by taking a limit of Eqn. (8)), we have
$G_{N,\theta}(\vec{x})=\int_{-\infty}^{\infty}{\frac{e^{-\frac{a^{2}}{2N}}}{\sqrt{2\pi N}}\cdot\exp\left(a\sum_{i=1}^{N}x_{i}\right)da}\\ =\exp\left(\frac{N}{2}p_{1}(\vec{x})^{2}\right)=\exp\left(\frac{N}{2}M_{(2)}(\vec{x})+NM_{(1,1)}(\vec{x})\right).$
One can check that $a(N)$ satisfies the fixed temperature LLN with $m_{k}=0$ for all $k$. This is revealed by the fact that $\lim_{N\rightarrow\infty}\frac{a^{N}_{(1,1)}}{N^{2}}=0$ and $a^{N}_{\lambda}=0$ for all $\lambda\neq(1,1)$. On the other hand, $G_{N,\theta}(\vec{x})$ does not satisfy the conditions proposed by [BCG22, Claim 9.1], which are given by monomials of partial derivatives.